1. Rank the following proposals in terms of their ability to increase the equilibrium level of income :
a- Government spending and taxes increase by x
b- Government spending and taxes decrease by x
c- Government spending increase by x
d- Taxes increase by x
2. Suppose that MPC = 0.9 and MP to import = 0.10 and an autonomous increase in Exports = $10 . Calculate :
a. The level of the multiplier
b. The change in the level of imports
( Hint : k = 1 / 1- MPC + MPM )
1
Expert's answer
2019-05-27T12:18:55-0400

1.

a- If government spending and taxes increase by x, then the equilibrium level of income can either increase or decrease.

b- If government spending and taxes decrease by x, then the equilibrium level of income can either increase or decrease.

c- If government spending increase by x, then the equilibrium level of income will increase.

d- If taxes increase by x, then the equilibrium level of income will decrease.

2. Suppose that MPC = 0.9, MPI = 0.10 and an autonomous increase in Exports = $10.

a. The level of the multiplier is:

k = 1/(1 - 0.9 + 0.1) = 1.25.

b. The change in the level of imports is:

10×1.25×0.1 = 1.25.